The wall of house, 7m wide and 6m high is made from 0.3m thick brick with kT =0.6 W/m-K. The surface 

temperature on the inside of the wall is 16˚C and that on the outside is 6˚C and that on the outside is 6˚C. Find

the total heat loss per second through it.


1
Expert's answer
2021-05-06T07:41:53-0400

"Q=k\frac{\Delta T}{\Delta x}At"


"A=7\cdot6=42\ (m^2)"

"t=1\ (s)"

"\Delta T=16-6=10\ (°C)"

"\Delta x=0.3\ (m)"

"k=0.6\ (W/(m\cdot K))"


"Q=0.6\cdot\frac{10}{0.3}\cdot42\cdot1=840\ (J)" . Answer